1. A solid conducting sphere of radius 40 cm has a charge of 5 µC. If the potential is zero

at infinity. Find the value of the potential at the following distances from the center of

the sphere:

a) 50 cm b) 40 cm c) 25 cm



1
Expert's answer
2021-06-17T09:30:41-0400

"\varphi= \begin{cases}
 \frac 32\frac{kq}{R}- \frac{kqr^2}{2R^3}&\text{if } r<R, \\
 \frac{kq}{r}&\text{if } r≥R;
\end{cases}"

a)

"\varphi=\frac{9\cdot 10^9\cdot 5\cdot 10^{-6}}{0.5}=90~kV,"

b)

"\varphi=\frac{9\cdot 10^9\cdot 5\cdot 10^{-6}}{0.4}=112.5~kV,"

c)

"\varphi=\frac 32\cdot \frac{9\cdot 10^9\cdot 5\cdot 10^{-6}}{0.4} -\frac{9\cdot 10^9\cdot 5\cdot 10^{-6}\cdot 0.25^2}{2\cdot 0.4^3}= 146.8~kV."